OpenStudy (anonymous):

2 STEP EQUATION 16x -19=113 - 6x

OpenStudy (anonymous):

First, get all your x's on one side and all your regular numbers on the other side of the equation. First, we'll add 6x + 19 to cancel out the -6x and -19. 16x - 19 = 113 - 6x +6x +19 +19 +6x ------------------ 22x = 132 Part one is done. Now we move onto part two: division. We'll divide both sides by 22 to get a lone x. 22x = 132 22x/22 = 132/22 Which will simplify to: x = 132/22 x = 6

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem! Glad I could help you :)